OpenStudy (anonymous):

integral(sqrt(4+x^2), x)

OpenStudy (anonymous):

Integrals with this form: \[ \int \sqrt{a^2+b^2x^2}dx \]Require the trig substitution \(x = \frac{a}{b}\tan(\theta)\). It's just an integration technique you have to memorize to be good at integrals. Integrals are tricky, because not all functions have an anti derivative.

OpenStudy (anonymous):

how to get the ln.????

OpenStudy (anonymous):

integral((sec(x))^3, x) how to solve it.. i stuck until this stage only.???